# A Multi-Fidelity Parametric Framework for Reduced-Order Modeling using Optimal Transport-based Interpolation: Applications to Diffused-Interface Two-Phase Flows

### 04 Mar 2026 v1

@@ 3. Multi-Fidelity Residual Interpolation
In many applications, one has access to models of varying fidelity, which could originate from different assumptions or discretization. A high-fidelity model provides accurate solutions ${\bm{u}}_{\mathrm{HF}}(t)$ but at substantial computational cost, while a low-fidelity solution ${\bm{u}}_{\mathrm{LF}}(t)$ can be evaluated cheaply but with reduced accuracy. In particular, the LF model might arise from a coarser spatial discretization, a larger time step, or simplified physics. The goal of multi-fidelity modeling is to combine the strengths of both: exploiting the efficiency of LF models to obtain many simulation data to complete the missing information obtained from a limited number of HF simulations [49]. In practice, this strategy allows for an efficient exploration of the solution manifold, both in the offline and in the online phases, exploiting a cheaper model to generate many snapshots and building a non-intrusive real-time surrogate, respectively.
A straightforward strategy is to model the residual or discrepancy field ${\bm{r}}(t)={\bm{u}}_{\mathrm{HF}}(t)-{\bm{u}}_{\mathrm{LF}}(t)$, which captures the difference between the two models. The HF solution can then be approximated as ${\bm{u}}_{\mathrm{HF}}(t)\approx{\bm{u}}_{\mathrm{LF}}(t)+{\bm{r}}_{\text{approx}}(t)$, where ${\bm{r}}_{\text{approx}}(t)$ is a reduced-order and efficient-to-evaluate representation of the residual. In standard multi-fidelity ROMs, one constructs a POD basis from residual snapshots and uses regression techniques to predict the reduced coefficients at new time instances. However, for problems with moving features, the residual field itself exhibits advective dynamics that are poorly captured by linear interpolation in the POD coefficient space, motivating our OT-based approach.
We propose to use displacement interpolation to model the temporal evolution of the residual field, developing the Multi-Fidelity OT-ROM (MF-OT-ROM) strategy illustrated in Figure 3. Given HF and LF snapshots at checkpoint times $\{t_{k}\}_{k=1}^{N_{c}}$, we compute residual checkpoints ${\bm{r}}_{k}={\bm{u}}_{\mathrm{HF}}(t_{k})-{\bm{u}}_{\mathrm{LF}}(t_{k})$ and apply OT-based interpolation to generate ${\bm{r}}_{\text{interp}}(t^{*})$ for any query time $t^{*}$. Since residual fields typically contain both positive and negative values, we apply a simple decomposition into positive and negative parts, ${\bm{r}}={\bm{r}}^{+}-{\bm{r}}^{-}$, with each part interpolated separately before recombination, so that the final multi-fidelity approximation is given by
@@
This multi-fidelity approach offers clear advantages over simply applying standard interpolation (e.g., linear or spline) to residual coefficients in a POD basis, especially in this context. Indeed, the OT-based interpolation inherently respects the geometric structure and movement of features within the residual field, and thus, for phase-field problems, where the residual is often concentrated at the moving interface, the method can effectively “transport” the correction from its position at $t_{k}$ to its new position at $t_{k+1}$, rather than just fading out the old correction and fading in the new one.
@@ 4. Parametric Displacement Interpolation
In the previous sections, we have specifically considered time-dependent problems as the natural framework where to exploit our strategy. Nevertheless, being a data-driven strategy, both temporal or parametric nature of the moving features can be treated in a similar way, thus we are interested in extending the strategy proposed in [38] to perform displacement interpolation for problems with parametric dependencies. More specifically, in the formulation presented so far, time $t$ serves as the interpolation coordinate in the solution manifold, with OT plans computed between consecutive temporal checkpoints, while we aim at including additional parameters $\mu\in\mathcal{P}\subset\mathbb{R}^{p}$, modeling physical characteristics of the systems, such as material properties and boundary conditions, or geometric properties.
In Figure 4, we present the procedure to exploit a dataset of dynamical systems corresponding to different parameter values performing a two-stage displacement interpolation strategy to obtain synthetic checkpoints for newly given parameters and recover their temporal evolution. Practically speaking, given solution snapshots at training parameter values $\{\mu_{j}\}_{j=1}^{N_{p}}$ and checkpoint times $\{t_{k}\}_{k=1}^{N_{c}}$, we can interpolate along both coordinate directions: temporally at fixed parameter, or parametrically at fixed time. For a test parameter $\mu^{*}\in(\mu_{j},\mu_{j+1})$, OT plans are computed between snapshots at adjacent training parameters for each checkpoint, and synthetic checkpoints are generated via parametric displacement interpolation with factor $\alpha_{\mu}=(\mu^{*}-\mu_{j})/(\mu_{j+1}-\mu_{j})$. Finally, temporal displacement interpolation is exploited to connect these and obtain efficient and reliable predictions for any query time. The extension to higher-dimensional parameter spaces can be done in several ways, exploiting additional knowledge on the parametric dependency, considering tensor product of grids, or using Wasserstein barycenters. Since all these extensions could potentially require a dramatic increase of the computational cost when the number of parameters increases, we postpone the investigation of an efficient strategy to following works.
@@
Finally, we introduce the Parametric Multi-Fidelity OT-ROM (PMF-OT-ROM), combining the multi-fidelity framework from Section 3 with the parametric setting. We start by noticing that two strategies are possible, depending on the field being considered. The first approach entails applying parametric displacement interpolation directly to the HF training data to generate synthetic HF checkpoints ${\bm{u}}_{\mathrm{HF}}^{\text{synth}}(t_{k};\mu^{*})$. A cheap LF simulation is run at $\mu^{*}$, and the synthetic residuals ${\bm{r}}^{\text{synth}}(t_{k})={\bm{u}}_{\mathrm{HF}}^{\text{synth}}(t_{k})-{\bm{u}}_{\mathrm{LF}}(t_{k})$ are interpolated in time to produce the final approximation ${\bm{u}}_{\text{approx}}(t)={\bm{u}}_{\mathrm{LF}}(t)+{\bm{r}}_{\text{interp}}(t)$.
On the other hand, the second strategy computes residuals at each training parameter value, then applies parametric displacement interpolation to these residual fields to generate synthetic residual checkpoints ${\bm{r}}^{\text{synth}}(t_{k};\mu^{*})$. Both strategies avoid constructing a monolithic ROM over the joint $(t,\mu)$ space, rather they explore the displacement interpolation along the two coordinates shifting the focus from the physical to the residual fields. The former approach is conceptually simpler, especially for applications where the high-fidelity solution has relevant physical meaning, while residual interpolation could be less intuitive, but preserving the same characteristics of moving features, it may be advantageous when its behavior with respect to $\mu$ varies more smoothly than the full solution.
Finally, we remark that a key assumption is that the solution (or the residual) evolves smoothly with $\mu$, thus problems exhibiting bifurcations [53, 27] or sharp transitions in parameter space [28, 59], more sophisticated strategies such as local model order reduction [4, 23] or adaptive sampling [50, 36] may be required.
@@ 5. Problem Definition: Five equation model
The proposed methodological extension of OT-based ROMs are applied in this work to the simulation of compressible two-phase flows. Specifically, the full-order model relies on the five-equation formulation originally introduced by Allaire et al. [3]. Compared to the original model, the present approach incorporates an additional transport equation for the level-set function $\psi$, which is employed to evaluate the interface normal vectors, following a strategy similar to that proposed by Al-Salami et al. [2].
The five-equation model reads:
$\frac{\partial\psi}{\partial t}+\textbf{u}\cdot\nabla\psi$ $=0,$ (20)
$\frac{\partial\phi_{1}}{\partial t}+\textbf{u}\cdot\nabla\phi_{1}$ $=\nabla\cdot\textbf{a}_{1},$ (21)
$\frac{\partial(\rho_{l}\phi_{l})}{\partial t}+\nabla\cdot(\rho_{l}\phi_{l}\textbf{u})$ $=\nabla\cdot\textbf{R}_{l},\quad l=\{1,2\},$ (22)
$\frac{\partial(\rho\textbf{u})}{\partial t}+\nabla\cdot(\rho\textbf{u}\otimes\textbf{u}+p\mathbb{I})$ $=\nabla\cdot(\textbf{f}\otimes\textbf{u})+\nabla\cdot\bm{\tau}+\sigma\kappa\nabla\phi_{1},$ (23)
$\frac{\partial(\rho E)}{\partial t}+\nabla\cdot[(\rho E+p)\textbf{u}]$ $=\nabla\cdot(\textbf{f}g)+\sum_{l=1}^{2}\nabla\cdot(\rho_{l}H_{l}\textbf{a}_{l})$
$\quad+\nabla\cdot(\bm{\tau}\cdot\textbf{u})+\sigma\kappa\nabla\phi_{1}\cdot\textbf{u},$ (24)
where $\phi_{l}$ denotes the $l^{\mathrm{th}}$ phase field and $H_{l}$ its specific enthalpy. The variable $\rho$ is the mixture density, $\rho\textbf{u}$ the total momentum, and $p$ the pressure. The total energy is defined as $\rho E=\rho e+\frac{1}{2}\rho\|\textbf{u}\|^{2}$, i.e., the sum of internal and kinetic energy. The parameter $\sigma$ denotes the surface tension coefficient and $\kappa=-\nabla\cdot\widehat{\textbf{n}}_{1}$ the interface curvature.
The following auxiliary quantities have been introduced:
@@
$\textbf{R}_{l}=\rho_{l}^{(0)}\textbf{a}_{l},\qquad\textbf{f}=\sum_{l=1}^{2}\textbf{R}_{l},\qquad g=\frac{1}{2}\|\textbf{u}\|^{2}.$ (26)
The parameters $\varepsilon$ and $\Gamma$ characterize the Allen-Cahn regularization. In particular, $\varepsilon$ controls the interface thickness and is typically chosen proportional to the grid size, while $\Gamma$ regulates the magnitude of the regularization relative to advection and is commonly scaled with the maximum velocity magnitude in the computational domain.
@@
$\bm{\tau}=2\mu\left(\mathbb{S}-\frac{1}{3}(\nabla\cdot\textbf{u})\mathbb{I}\right),$
where $\mu$ is the mixture dynamic viscosity, evaluated as $\mu=\mu_{1}\phi_{1}+\mu_{2}\phi_{2}$, and $\mathbb{S}=[\nabla\textbf{u}+(\nabla\textbf{u})^{\intercal}]/2$ denotes the strain-rate tensor.
@@
In the following section, we introduce the numerical benchmarks employed to assess the OT-based methodologies. The spatial discretization of the five-equation model, in both kinetic and fully coupled configurations, is performed using a high-order Spectral Difference scheme (see [72] for further details regarding the numerical discretization of the five equation model).
In kinetic tests, the phase field evolves under a prescribed analytical velocity field, whereas in the fully coupled configuration the complete system of governing equations is solved. As representative kinetic benchmark, we consider the Rider-Kothe vortex, while the Rayleigh-Taylor instability is adopted as the fully coupled test case. Both problems are inherently unsteady and involve explicit parametric dependencies in their initial conditions.
In all cases, the analysis focuses on the accurate prediction of interfacial dynamics. The primary quantity of interest is the phase-field variable $\phi_{1}$. Although the same framework can be extended to other flow variables with different physical interpretations, $\phi_{1}$ is selected due to its central role in the modeling of two-phase compressible flows.
@@ 6. Numerical Results
In this section, we assess the performance of the proposed MF-OT-ROM and PMF-OT-ROM frameworks on benchmark problems based on the five equation model described in Section 5. In order to present a comprehensive evaluation of the performance of both frameworks, for the MF-OT-ROM, we demonstrate the accuracy of the residual interpolation and the improvement over the standalone LF model, while for the PMF-OT-ROM, we evaluate the model’s ability to predict solutions for unseen parameter values.

### 27 Mar 2026 v2

@@ 4. Multi-Fidelity Residual Interpolation
In many applications, one has access to models of varying fidelity, which could originate from different assumptions or discretization. A high-fidelity model provides accurate solutions ${\bm{u}}_{\mathrm{HF}}(t)$ but at substantial computational cost, while a low-fidelity solution ${\bm{u}}_{\mathrm{LF}}(t)$ can be evaluated cheaply but with reduced accuracy. In particular, the LF model might arise from a coarser spatial discretization (the setting considered in this work), a larger time step, or simplified physics. The goal of multi-fidelity modeling is to combine the strengths of both: exploiting the efficiency of LF models to obtain many simulation data to complete the missing information obtained from a limited number of HF simulations [62, 19, 47, 58]. In practice, this strategy allows for an efficient exploration of the solution manifold, both in the offline and in the online phases, exploiting a cheaper model to generate many snapshots and building a non-intrusive real-time surrogate, respectively.
A straightforward strategy is to model the residual or discrepancy field ${\bm{r}}(t)={\bm{u}}_{\mathrm{HF}}(t)-{\bm{u}}_{\mathrm{LF}}(t)$, which captures the difference between the two models. The HF solution can then be approximated as ${\bm{u}}_{\mathrm{HF}}(t)\approx{\bm{u}}_{\mathrm{LF}}(t)+{\bm{r}}_{\text{approx}}(t)$, where ${\bm{r}}_{\text{approx}}(t)$ is a reduced-order and efficient-to-evaluate representation of the residual, that, for example, can be constructed via a POD-regression strategy on the residual snapshots. However, for problems with moving features, the residual field itself exhibits advective dynamics that are poorly captured by a linear combination in the POD coefficient space, motivating our OT-based approach.
Thus, we propose to use displacement interpolation to model the temporal evolution of the residual field, developing the Multi-Fidelity OT-ROM (MF-OT-ROM) strategy illustrated in Figure 3.
@@
Given HF and LF snapshots at checkpoint times $\{t_{k}\}_{k=1}^{N_{c}}$, we compute residual checkpoints ${\bm{r}}_{k}={\bm{u}}_{\mathrm{HF}}(t_{k})-{\bm{u}}_{\mathrm{LF}}(t_{k})$. In practice, for resolution-based multi-fidelity approaches, the LF and HF models are typically defined on different spatial grids. In all the numerical experiments considered here, we employ Cartesian grids with refinement factors of $2$ or $4$, therefore, the LF solution is interpolated onto the HF grid so that the residual can be evaluated on a common spatial discretization. We then apply OT-based interpolation to generate ${\bm{r}}_{\text{interp}}(t^{*})$ for any query time $t^{*}$. Since residual fields typically contain both positive and negative values, we apply a simple decomposition into positive and negative parts, ${\bm{r}}={\bm{r}}^{+}-{\bm{r}}^{-}$, with each part interpolated separately before recombination, so that the final multi-fidelity approximation is given by
This multi-fidelity approach offers clear advantages over simply applying standard interpolation (e.g., linear or spline) to residual coefficients in a POD basis, especially in this context. Indeed, the OT-based interpolation inherently respects the geometric structure and movement of features within the residual field, and thus, for diffused-interface models, where the residual is often concentrated at the moving interface, the method can effectively “transport” the correction from its position at $t_{k}$ to its new position at $t_{k+1}$, rather than just fading out the old correction and fading in the new one.
@@ 5. Parametric Displacement Interpolation
In the previous sections, we have specifically considered time-dependent problems as the natural framework where to exploit our strategy. Nevertheless, being a data-driven strategy, both temporal and parametric nature of the moving features can be treated in a similar way, thus we are interested in extending the strategy proposed in [48] to perform displacement interpolation for problems with parametric dependencies. More specifically, in the formulation presented so far, time $t$ serves as the interpolation coordinate in the solution manifold, with OT plans computed between consecutive temporal checkpoints, while we aim at including additional parameters $\mu\in\mathcal{P}\subset\mathbb{R}^{p}$, modeling physical characteristics of the systems, such as material properties and boundary conditions, or geometric properties.
In Figure 4, we present the procedure to exploit a dataset of dynamical systems corresponding to different parameter values performing a two-stage displacement interpolation strategy to obtain synthetic checkpoints for newly given parameters and recover their temporal evolution. Practically speaking, given solution snapshots at scalar ($p=1$) training parameter values $\{\mu_{j}\}_{j=1}^{N_{p}}$ and checkpoint times $\{t_{k}\}_{k=1}^{N_{c}}$, we can interpolate along both coordinate directions: temporally at fixed parameter, or parametrically at fixed time. Following this latter option, for a test parameter $\mu^{*}\in(\mu_{j},\mu_{j+1})$, OT plans are computed between snapshots at adjacent training parameters for each checkpoint, and synthetic snaphshots are generated via parametric displacement interpolation with “virtual parameter” $\alpha_{\mu}=(\mu^{*}-\mu_{j})/(\mu_{j+1}-\mu_{j})$. Next, temporal displacement interpolation is exploited to connect these generated checkpoints, and obtain efficient and reliable predictions for any query time. We remark that the extension to higher-dimensional parameter spaces can be done in several ways, exploiting additional knowledge on the parametric dependency, considering tensor product of grids. Since all these extensions could potentially require a dramatic increase of the computational cost when the number of parameters increases, we postpone the investigation of an efficient strategy to following works.
@@
By combining the multi-fidelity framework from Section 4 with the parametric setting, here we also aim at introducing the Parametric Multi-Fidelity OT-ROM (PMF-OT-ROM). We start by noticing that two strategies are possible, depending on the field being considered and the characteristics of the problem at hand. The first approach entails applying parametric displacement interpolation directly to the HF training snapshots data to generate synthetic HF checkpoints ${\bm{u}}_{\mathrm{HF}}^{\text{synth}}(t_{k};\mu^{*})$. A cheap LF simulation is run at $\mu^{*}$, and the synthetic residuals ${\bm{r}}^{\text{synth}}(t_{k};\mu^{*})={\bm{u}}_{\mathrm{HF}}^{\text{synth}}(t_{k};\mu^{*})-{\bm{u}}_{\mathrm{LF}}(t_{k};\mu^{*})$ are interpolated in time to produce the final approximation ${\bm{u}}_{\text{approx}}(t;\mu^{*})={\bm{u}}_{\mathrm{LF}}(t;\mu^{*})+{\bm{r}}_{\text{interp}}(t;\mu^{*})$. On the other hand, one could compute residuals at each training parameter value ${\bm{r}}^{\text{synth}}(t_{k};\mu_{j})={\bm{u}}_{\mathrm{HF}}^{\text{synth}}(t_{k};\mu_{j})-{\bm{u}}_{\mathrm{LF}}(t_{k};\mu_{j})$, apply parametric displacement interpolation to these residual fields to generate synthetic residual checkpoints ${\bm{r}}^{\text{synth}}(t_{k};\mu^{*})$, and finally perform temporal displacement interpolation to have a continuous corrected model in time.
Both strategies avoid constructing a monolithic ROM over the joint $(t,\mu)$ space, rather they explore the displacement interpolation along the two coordinates shifting the focus from the physical to the residual fields. The former approach is conceptually simpler, especially for applications where the high-fidelity solution has relevant physical meaning, while residual interpolation could be less intuitive, but preserving the same characteristics of moving features, it may be advantageous when its behavior with respect to $\mu$ varies more smoothly than the full solution.
In terms of computational cost, it is important to highlight two advantages of the former approach (i.e., interpolating HF solutions first and then computing the residual). First, the DI step needs to be performed only three times instead of four, since the phase-field is strictly non-negative and therefore does not require a prior decomposition into positive and negative parts. Second, this approach only requires HF solutions along the neighboring parametric trajectories, while for the unseen parameter $\mu^{*}$ only the LF solution is needed. Instead, interpolating residuals in both time and parameter space requires LF solutions also for the neighboring trajectories. These additional computations ar thus essential to reconstruct the residual associated with the trajectory corresponding to $\mu^{*}$. For these reasons, in the numerical experiments considered in this work we adopt this approach within the parametric multi-fidelity strategy.
A key assumption underlying the proposed approach is that the solution (or the residual field) varies smoothly with respect to the parameter $\mu$. In problems where the parametric dependence exhibits bifurcations [66, 32] or sharp transitions in parameter space [33, 73], the interpolation-based strategy may become less effective. In such situations, more sophisticated techniques such as local model order reduction [4, 26] or adaptive sampling strategies [63, 44] may be required.
@@ 6. Problem Definition: Five equation model
The proposed methodological extensions of OT-based ROMs are applied in this work to the numerical simulation of compressible two-phase flows using diffuse-interface methods. Specifically, the full-order model relies on the five-equation formulation originally introduced by Allaire et al. [3]. From the Baer-Nunziato seven-equation model [7], assuming mechanical and thermodynamic equilibrium between the two phases, the following five-equation model can be derived:
$\frac{\partial\psi}{\partial t}+\textbf{v}\cdot\nabla\psi$ $=0,$ (19)
$\frac{\partial\phi_{1}}{\partial t}+\textbf{v}\cdot\nabla\phi_{1}$ $=\nabla\cdot\textbf{a}_{1},$ (20)
$\frac{\partial(\rho_{l}\phi_{l})}{\partial t}+\nabla\cdot(\rho_{l}\phi_{l}\textbf{v})$ $=\nabla\cdot\textbf{R}_{l},\quad l=\{1,2\},$ (21)
$\frac{\partial(\rho\textbf{v})}{\partial t}+\nabla\cdot(\rho\textbf{v}\otimes\textbf{v}+p\mathbb{I})$ $=\nabla\cdot(\textbf{f}\otimes\textbf{v})+\nabla\cdot\bm{\tau}+\sigma\kappa\nabla\phi_{1},$ (22)
$\frac{\partial(\rho E)}{\partial t}+\nabla\cdot[(\rho E+p)\textbf{v}]$ $=\nabla\cdot(\textbf{f}g)+\sum_{l=1}^{2}\nabla\cdot(\rho_{l}H_{l}\textbf{a}_{l})+\nabla\cdot(\bm{\tau}\cdot\textbf{v})+\sigma\kappa\nabla\phi_{1}\cdot\textbf{v},$ (23)
consisting of one evolution equation for the phase-field (20), where $\phi_{1}$ denotes the phase-field referred to the first phase, two separate mass conservation laws for the individual phases (21), one balance equation for the mixture total momentum (22) and one balance equation for the total energy of the mixture (23) with $H_{l}$ the specific enthalpy of the $l^{\mathrm{th}}$ phase. Compared to the original model, the present approach incorporates an additional transport equation for the level-set function $\psi$ (19), which is used to evaluate the interface normal vectors, following a strategy similar to that proposed by Al-Salami et al. [2]. The variable $\rho$ is the mixture density, $\rho\textbf{v}$ the total momentum, and $p$ the pressure. The total energy is defined as $\rho E=\rho e+\frac{1}{2}\rho\|\textbf{v}\|^{2}$, i.e., the sum of internal and kinetic energy. The parameter $\sigma$ denotes the surface tension coefficient, $\kappa=-\nabla\cdot\widehat{\textbf{n}}_{1}$ the interface curvature, and the following auxiliary quantities are commonly introduced:
@@
$\textbf{R}_{l}=\rho_{l}^{(0)}\textbf{a}_{l},\qquad\textbf{f}=\sum_{l=1}^{2}\textbf{R}_{l},\qquad g=\frac{1}{2}\|\textbf{v}\|^{2}.$ (25)
The regularization term $\mathbf{a}_{1}$ in Equation (20) is related to the conservative Allen-Cahn formulation proposed by [16], and is governed by two main parameters, $\varepsilon$ and $\Gamma$. The former controls the interface thickness and is typically chosen to be proportional to the grid size, while the latter regulates the magnitude of the regularization relative to advection and is commonly scaled with the maximum velocity magnitude in the computational domain. All the additional terms involving contributions that depend directly on $\mathbf{a}_{l}$ are classically added to the remaining conservation laws in order to ensure consistency and preserve kinetic energy and entropy in the overall coupled system [42].
@@
$\bm{\tau}=2\mu\left(\mathbb{S}-\frac{1}{3}(\nabla\cdot\textbf{v})\mathbb{I}\right),$
where $\mu$ is the mixture dynamic viscosity, evaluated as $\mu=\mu_{1}\phi_{1}+\mu_{2}\phi_{2}$, and $\mathbb{S}=[\nabla\textbf{v}+(\nabla\textbf{v})^{\intercal}]/2$ denotes the strain-rate tensor.
@@
The spatial discretization of the five-equation model, in both kinetic and fully coupled configurations, is performed using a high-order Spectral Difference scheme (see [86] for further details regarding the numerical discretization of the five-equation model).
In kinetic tests, the phase-field evolves under a prescribed analytical velocity field, whereas in the fully coupled configuration the complete system of governing equations is solved. As representative kinetic benchmark, we consider the Rider-Kothe vortex [72], while the Rayleigh-Taylor instability [87] is adopted as the fully coupled test case. Both problems are inherently unsteady and involve explicit parametric dependencies in their initial conditions.
In all cases, the analysis focuses on the accurate prediction of interfacial dynamics. The primary quantity of interest is consequently represented by the phase-field variable $\phi_{1}$. Although the same framework can be extended to other flow variables with different physical interpretations, $\phi_{1}$ is selected due to its central role in the modeling of two-phase compressible flows. Furthermore, for the sake of simplicity, we will refer to $\phi_{1}$ simply as $\phi$ for the remainder of the paper.
@@ 7. Numerical Results
In this section, we assess the performance of the proposed MF-OT-ROM and PMF-OT-ROM frameworks on benchmark problems based on the diffuse-interface models described in Section 6. In order to present a comprehensive evaluation of the performance of both frameworks, for the MF-OT-ROM, we demonstrate the accuracy of the corrected residual interpolation and the improvement over the standalone uncorrected LF solution for single trajectories involving multiple fidelities, while for the PMF-OT-ROM, we evaluate the model’s ability to improve LF solutions for completely unseen parametric trajectories.